OpenStudy (anonymous):

find four consecutive even numbers whose sum is 156

OpenStudy (lgbasallote):

let x = even number x+2 = consecutive even number x+4 = next x+6 = last consecutive number x + x + 2 + x + 4 + x + 6 = 156 4x +12 = 156 4x = 156-12 4x = 144 x = 36 x+ 2 = 38 x+4 = 40 x+6 = 42